OpenStudy (anonymous):

factor trinomial by grouping r^2-3rs-40s^2

OpenStudy (anonymous):

\[r^2-3rs-40s^2\] \[r^2-2\frac{3}{2}rs+\frac{9}{4}s^2-\frac{9}{4}s^2-40s^2\] \[(r-\frac{3s}{2})^2-\frac{169}{4}s^2\] \[(r-\frac{3s}{2}+\frac{13}{2}s)(r-\frac{3s}{2}-\frac{13}{2}s)\] \[(r+\frac{10s}{2})(r-\frac{16s}{2})\]

OpenStudy (anonymous):

wow, thanks so much

OpenStudy (anonymous):

you're welcome!